OpenStudy (anonymous):

integrate 1/1+x^2 dx

OpenStudy (anonymous):

arctan(x) + C

OpenStudy (anonymous):

that i know.. how u got there is what i want to know

OpenStudy (anonymous):

let y = arctan(x) then x = tan(y) by implicit differentiation, sec^2(y) * dy/dx = 1 thus dy/dx = cos^2(y) since x = tan(y), then cos(y) = 1/sqrt(1+x^2) thus cos^2(y) = 1/(1 + x^2) thus dy/dx = 1/(1 + x^2) hence the derivative of arctan(x) is 1/(1 + x^2) hence the integral of 1/(1 + x^2) is arctan(x)

OpenStudy (anonymous):

You could do it with trigonometric substitution. But if you haven't covered that then james^^ has about the best explanation.

OpenStudy (anonymous):

The basic trig inverse functions, sin, cos, tan, etc have a certain form. You need to become familiar with these and be able to recognize them.